# Assessing differences between local galaxy dust attenuation and point source extinction within the same environments

### v1

@@ 3.2.2 Local Observations
We can examine how well these results line up with observations by examining some preliminary dust attenuation values obtained for the All-weather MUse Supernova Integral Field Nearby Galaxies (AMUSING; Galbany et al., 2016) data for the NGC 1448 galaxy. This integral field spectroscopy (IFS) data was matched to GALEX (Martin et al., 2005), HST, 2MASS (Skrutskie et al., 2006) and WISE (Wright et al., 2010) photometry using the piXedfit Python package (Abdurro’uf et al., 2021). The Prospector Python package (Johnson et al., 2021) was used to fit each pixel in a subsection of NGC 1448, with the attenuation properties being expressed in terms of the effective optical depth $\tau_{V_{eff}}$ and the dust index $n$, which controls the slope of the Kriek & Conroy (2013) attenuation curve.
The results for the fitted $\tau_{V_{eff}}$-$n$ trends are displayed in Fig. 14. While these fits are still preliminary, it appears that the same trend observed in Fig. 10, 12 and 13 is present in the AMUSING data, even if expressed through different parameters. It is not totally clear whether the $n$ values are converging toward a value representative of the extinction properties of the medium, or if they are simply clustering near the edge of the prior distribution. Furthermore, our simulations and fits match the results obtained by Decleir et al. (2019), who find a similar increase of the attenuation curve slope with $A_{V_{eff}}$ for resolved observations of NGC 628. These results give some confidence to the applicability of the inferred simulated trends to real galaxies.
Figure 14: Dust index $n$ as a function of the optical depth $\tau_{V_{eff}}$ for fits of a subsection of NGC 1448, using AMUSING IFS data matched with GALEX, HST, 2MASS and WISE photometry. Values for individual pixels are shown in blue, while binned median values are shown in orange.
@@ 4.1 Dust Attenuation and Stellar Mass
Among others, Garn & Best (2010), Zahid et al. (2013), Salim et al. (2018), Galbany et al. (2022) and Duarte et al. (2023) show that, for star-forming galaxies, there is an increase of $A_{V_{eff}}$ with the stellar mass. It is fair to assume that this is related to an increase in the dust density in high-mass galaxies. Salim et al. (2018) and Duarte et al. (2023) also show that high-mass star-forming galaxies exhibit larger values of $R_{V_{eff}}$. From Fig. 6 it is clear that an increase in the dust density in a galaxy leads directly to an increase in the median values of both $A_{V_{eff}}$ and $R_{V_{eff}}$. Assuming that there is indeed a link between the mass of a galaxy and its dust density, our simulations are able to fully explain the observed increase of $R_{V_{eff}}$ with stellar mass. Crucially, they do so without the need to consider the differing dust types or source SEDs.
@@
Another important result discussed in Zahid et al. (2013), Salim et al. (2018) and Duarte et al. (2023) is that elliptical galaxies are, in general, expected to have much lower values of $A_{V_{eff}}$ and $R_{V_{eff}}$ than high-mass star-forming galaxies. It is this fact that makes it possible for the entire high-mass population to actually have a lower median $R_{V_{eff}}$ than the one observed for the low-mass galaxies. However, this is something our simulations cannot reproduce. If one assumes that the median $A_{V_{eff}}$ observed for a population of elliptical galaxies should be roughly the same as the one observed for a population of low-mass star-forming galaxies, which is corroborated by Zahid et al. (2013), Salim et al. (2018) and Duarte et al. (2023), we can use this assumption to properly normalize the dust contents of both galaxy types. However, under these conditions, we invariably find that the ellipticals boast median $R_{V_{eff}}$ values higher than even those of the high-mass star-forming galaxies. There are two likely explanations. The first is that the simplistic geometries assumed in this work are not able to properly model these two types of galaxies in relation to one another. The second is that there is some fundamental difference between the two galaxy types that goes beyond geometry, such as a different dust composition with different optical properties.
@@
As suggested in Figs. 10, 12 and 13 and tentatively confirmed by Fig. 14, the analysis of full-galaxy field data for the regions of a face-on galaxy with higher $A_{V_{eff}}$ might offer an observational hint toward the true extinction curve slope $R_{V}$ of a given dust sample. While the robustness of this methodology in determining accurate $R_{V}$ values for real galactic dust mediums cannot be ascertained from our simple simulations, observational results are promising, as they show a similar increase in $R_{V_{eff}}$ with $A_{V_{eff}}$ as the one recovered for the simulations. Thus, this might be a very apt method to distinguish between different dust types, perhaps even within the same galaxy. As such, it might be able to provide a reliable way to relate galaxy attenuation data with the extinction of their hosted point sources, such as SNe. This would be invaluable in an effort to better constrain SN Ia calibration and the respective color-luminosity correction.
@@ 4.2 Source SED Impact
As mentioned above, when it comes to dust, the results presented in this paper are largely independent of the SED of the simulated sources, given that our simulations output precise fluxes for the emitted and observed fluxes. Because of this, we can avoid many of the uncertainties and degeneracies that are inherent in observations and SED fitting. However, there are a few points to be highlighted.
@@
One of the reasons why our results can be said to be SED independent is that, for any given model simulation, a single SED is used. This means that any effect of the source is factored out when the effective dust properties are computed. We do not, however, expect real galaxies to exhibit a SED that is uniform across all sources, i.e., we do not expect stellar populations of the same age everywhere. In this scenario, both $A_{V_{eff}}$ and $R_{V_{eff}}$ are also expected to be affected by the interplay of the various SEDs, particularly if sources of very different ages are not uniformly distributed within the galaxy and are subject to different levels of attenuation. This will contribute to further deviations from the medium extinction values. Likewise, while the overall luminosity of a galaxy is not a relevant component in our results, given that only the source distribution affects the simulated dust properties, it might play a role on the observed dust properties. In particular, if any of the flux components become dimmer than the limiting magnitude of the observational setup in question, we might end up recovering dust properties that are biased toward the most luminous regions.
@@ 5 Conclusions
In this work we have used radiative transfer simulations to explore how SN Ia dust extinction and host galaxy dust attenuation relate to one another. Our main findings are the following:
@@
    SN Ia extinction (or “point source extinction” as used throughout this paper) and host galaxy attenuation are fundamentally different phenomena. When observing a galaxy, the light scattered into the line of sight becomes a very relevant component of the total observed flux, lowering the value of $R_{V_{eff}}$. In addition, the fact that sources at different depths are contained within the same line of sight contributes to an increase in $R_{V_{eff}}$. When taken in conjunction, these two effects mean that the $R_{V_{eff}}$ measured for any given galaxy is fundamentally different from the optical extinction $R_{V}$ of its particular dust medium, even if small resolutions of $25$ pc are considered.
@@
    Global attenuation properties are primarily driven by observation angle. The global $A_{V_{eff}}-R_{V_{eff}}$ trends for observed galaxies can be qualitatively reproduced by a single simulated galaxy observed from different angles. Dust density and galaxy type serve simply to increase the scatter of this trend. We conclude that geometry, rather than dust density or composition, is the primary factor driving global attenuation.
@@
    An increase in galaxy dust density results in an increase of $R_{V_{eff}}$ across all observation angles. As such, assuming a correlation between stellar mass and dust density, the geometry of the observations is also able to explain the increase of $R_{V_{eff}}$ values with stellar mass for observed star-forming spiral galaxies.
@@
    When spiral and elliptical galaxies are analyzed together, our simulations fail to explain some of the observed behaviors. The literature suggests that ellipticals have $A_{V_{eff}}$ values similar to low-mass spirals. However, if this is assumed in our model, the recovered $R_{V_{eff}}$ values, which are also supposed to be in line with those of low-mass spirals, are much higher than expected. In particular, when taking all galaxies into account, we find that the high-mass population boasts a higher median $R_{V_{eff}}$ than its low-mass counterpart, which is the opposite of the observed trend. This could indicate either a failure in the modeling of one or both galaxy types or that there is something intrinsically different between the two galaxy types beyond geometry, namely their dust properties. In this regard, part of the origin of the mass-step correction in SN Ia distances could indeed be related to differing dust properties among galaxies.
@@
    The use of the full-galaxy field data opens up the possibility to quantifiably link host galaxy attenuation with point source (such as SNe Ia) extinction. While it is difficult to extract accurate information from individual locations, an analysis of the $A_{V_{eff}}-R_{V_{eff}}$ trend for a galaxy could allow one to infer an approximate value for the extinction $R_{V}$ of the medium and, most importantly, to distinguish between different dust types/compositions. This is tentatively confirmed by preliminary observed galaxy fits. Our examples suggest that, to recover $R_{V_{eff}}$ values within a $10\%$ error of the intrinsic extinction $R_{V}$, areas with high $A_{V_{eff}}$ should be considered, in the case of face-on galaxies. For edge-on spirals, however, the median $R_{V_{eff}}$ for the low $A_{V_{eff}}$ regions – although with a lot of scatter – appears to be better at tracing the intrinsic extinction $R_{V}$. In any case, it is crucial to ensure that these regions have properties that allow for the elimination of both the scattered flux component and the direct component coming from deeper regions of the galaxy.
@@ Acknowledgements.
J. D., S. G. and A. M. acknowledge support by FCT for CENTRA through the Project No. UIDB/00099/2020. J. D. also acknowledges support by FCT under the PhD grant 2023.01333.BD. This work was funded by ANID, Millennium Science Initiative, ICN12_009. Based in part on observations collected at the European Southern Observatory under ESO programme(s) 097.D-0408(A). Based in part on observations made with the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555, associated with programs #17070 and #17191. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. This work used public data from the GALEX Survey. GALEX is operated for NASA by the California Institute of Technology under NASA contract NAS5-98034. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration.

### v2

@@ 3.2.2 Local observations
We can examine how well these results line up with observations by examining some preliminary dust attenuation values obtained for the All-weather MUse Supernova Integral Field Nearby Galaxies (AMUSING; Galbany et al., 2016) data for the NGC 1448 galaxy. This integral field spectroscopy (IFS) data was matched to GALEX (Martin et al., 2005), HST, 2MASS (Skrutskie et al., 2006) and WISE (Wright et al., 2010) photometry using the piXedfit Python package (Abdurro’uf et al., 2021). The Prospector Python package (Johnson et al., 2021) was used to fit each pixel in a subsection of NGC 1448, with the attenuation properties being expressed in terms of the effective optical depth, $\tau_{V_{eff}}$, and the dust index, $n$, which controls the slope of the attenuation curve in Kriek & Conroy (2013), such that, on a rough approximation, we have $R_{V_{eff}}\sim\left(\frac{\lambda_{V}}{\lambda_{B}}\right)^{n}$. As was mentioned in Section 3.2.1, there appears to exist a bias in Prospector that results in the overestimation of $\tau_{V_{eff}}$. This does not pose a problem to our analysis, as we are primarily interested in the $\tau_{V_{eff}}$-$n$ trends.
The results for the fitted $\tau_{V_{eff}}$-$n$ trends are displayed in Fig. 14. While these fits are still preliminary, it appears that the same trend observed in Fig. 10, 12 and 13 is present in the AMUSING data, even if expressed through different parameters. It is not totally clear whether the $n$ values are converging toward a value representative of the extinction properties of the medium, or if they are simply clustering near the edge of the prior distribution. Furthermore, our simulations and fits match the results obtained by Decleir et al. (2019), who find a similar increase in the attenuation curve slope with $A_{V_{eff}}$ for resolved observations of NGC 628. These results give some confidence to the applicability of the inferred simulated trends to real galaxies.
Figure 14: Dust index, $n$, as a function of the optical depth $\tau_{V_{eff}}$ for fits of a subsection of NGC 1448, using AMUSING IFS data matched with GALEX, HST, 2MASS and WISE photometry. Values for individual pixels are shown in blue, while binned median values are shown in orange.
@@ 4.1 Dust attenuation and stellar mass
Among others, Garn & Best (2010), Zahid et al. (2013), Salim et al. (2018), Galbany et al. (2022) and Duarte et al. (2023) show that, for star-forming galaxies, there is an increase in $A_{V_{eff}}$ with stellar mass. It is fair to assume that this is related to an increase in the dust density in high-mass galaxies. Salim et al. (2018) and Duarte et al. (2023) also show that high-mass star-forming galaxies exhibit larger values of $R_{V_{eff}}$. From Fig. 6 it is clear that an increase in the dust density in a galaxy leads directly to an increase in the median values of both $A_{V_{eff}}$ and $R_{V_{eff}}$. Assuming that there is indeed a link between the mass of a galaxy and its dust density, our simulations are able to fully explain the observed increase in $R_{V_{eff}}$ with stellar mass. Crucially, they do so without the need to consider the differing dust types or source SEDs.
@@
Another important result discussed in Zahid et al. (2013), Salim et al. (2018) and Duarte et al. (2023) is that elliptical galaxies are, in general, expected to have much lower values of $A_{V_{eff}}$ and $R_{V_{eff}}$ than high-mass star-forming galaxies. It is this fact that makes it possible for the entire high-mass population to actually have a lower median $R_{V_{eff}}$ than the one observed for the low-mass galaxies. However, this is something our simulations cannot reproduce. If one assumes that the median $A_{V_{eff}}$ observed for a population of elliptical galaxies should be roughly the same as the one observed for a population of low-mass star-forming galaxies, which is corroborated by Zahid et al. (2013), Salim et al. (2018) and Duarte et al. (2023), we can use this assumption to properly normalize the dust contents of both galaxy types. However, under these conditions, we invariably find that the ellipticals have median $R_{V_{eff}}$ values higher than even those of the high-mass star-forming galaxies. There are two likely explanations. The first is that the simplistic geometries assumed in this work are not able to properly model these two types of galaxies in relation to one another. The second is that there is some fundamental difference between the two galaxy types that goes beyond geometry, such as a different dust composition with different optical properties, which makes it so that ellipticals have intrinsically lower $R_{V}$ values. This would support the claim that multiple dust properties can be present in SN Ia cosmological analyses, as has been argued by, among others, Brout & Scolnic (2021), González-Gaitán et al. (2021), and Wiseman et al. (2022).
@@
As is suggested in Figs. 10, 12, and 13 and tentatively confirmed by Fig. 14, the analysis of full-galaxy field data for the regions of a face-on galaxy with higher $A_{V_{eff}}$ might offer an observational hint at the true extinction curve slope $R_{V}$ of a given dust sample. While the robustness of this methodology in determining accurate $R_{V}$ values for real galactic dust mediums cannot be ascertained from our simulations, observational results are promising, as they show a similar increase in $R_{V_{eff}}$ with $A_{V_{eff}}$ as the one recovered for the simulations. Thus, the analysis of the local dust properties in highly attenuated regions of face-on galaxies might be a very apt method to distinguish between different dust types, perhaps even within the same galaxy. As such, it might be able to provide a reliable way to relate galaxy attenuation data with the extinction of their hosted point sources, such as SNe. This would be invaluable in an effort to better constrain SN Ia standardization and the respective color-luminosity correction.
@@ 4.2 Source SED impact
As was mentioned above, when it comes to dust, the results presented in this paper are largely independent of the SED of the simulated sources, given that our simulations output precise fluxes for the emitted and observed fluxes. Because of this, we can avoid many of the uncertainties and degeneracies that are inherent in observations and SED fitting. As an example, as was stated in Section 3.1, we recover the exact same values of $A_{V_{eff}}$ and $R_{V_{eff}}$ when using light source SEDs with either $t_{age}=10$ Myr or $t_{age}=5$ Gyr. This happens because the general scattering and attenuation effects found in this study are experienced by all sources of light, irrespective of their SED. However, there are a few points to be highlighted.
@@
One of the reasons why our results can be said to be SED independent is that, for any given model simulation, a single SED is used. This means that any effect of the source is factored out when the effective dust properties are computed. We do not, however, expect real galaxies to exhibit a SED that is uniform across all sources, i.e., we do not expect stellar populations of the same age everywhere. In this scenario, both $A_{V_{eff}}$ and $R_{V_{eff}}$ are also expected to be affected by the interplay of the various SEDs, particularly if sources of very different ages are not uniformly distributed within the galaxy and are subject to different levels of attenuation. We examine two specific examples to further illustrate this point.
We first look at a simulated spiral galaxy similar to the one described in Section 2.2, but with disk age $t_{disk}=100$ Myr and bulge age $t_{bulge}=5$ Gyr. Although not totally realistic, this simulation helps illustrate the impact of multiple SEDs differently distributed on the recovered effective dust properties. The 2D face-on and edge-on views for this galaxy are similar to those obtained for a spiral galaxy with identical morphology and a single SED (see Figs. 10 and 11). The main exception is that, in the edge-on view of the two-population simulation, $R_{V_{eff}}$ diverges much more from $R_{V_{Zubko}}$ as $A_{V_{eff}}$ increases.
Regarding the variation in the global dust properties across various observing angles, plotted in Fig. 15, we find that it substantially differs from the results displayed in Fig. 5. The presence of different star populations contributes to further complicate the geometric effects on the observed light, making it even harder to establish a link between extinction and attenuation. The fact that observed galaxy dust data are closer to the single SED simulation might indicate that the differences in age distributions between populations of the same galaxy are not as large as assumed in this extreme example.
Figure 15: Values of $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ for a simulated spiral galaxy with disk age $t_{disk}=100$ Myr and bulge age $t_{bulge}=5$ Gyr. Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). An aperture radius of $20$ kpc was used. The galaxy dust density has been normalized so that the total optical depth along the R axis is $\tau_{V_{R}}=5$. The dashed black line reflects the true extinction $R_{V_{Zubko}}$.
For a more realistic example, we examined a simulation based on the M51 galaxy model from the DustPedia [^5] http://dustpedia.astro.noa.gr/ archive and uses observations of M51 to define both the stellar and dust morphologies. Crucially, the model contains both old and young stars, with a much more realistic distribution than the one used for the previous example. The dust medium is characterized by a THEMIS dust mix (Jones, A. P. et al., 2017), with dust mass $M_{dust}=10^{8}M_{\odot}$. The global dust properties across various observing angles for the simulated M51 galaxy are plotted in Fig. 16. The results for this more realistic two-population simulation are much more in line with the ones recovered for both the single-population simulations and observations. This means that, for real galaxies, while the mixing of different SEDs might influence the individual values of $R_{V_{eff}}$, it will most likely not have a big impact on the overall shape of the $A_{V_{eff}}-R_{V_{eff}}$ trend.
Figure 16: Values of $R_{V_{eff}}$ as a function of $A_{V_{eff}}$ for a simulation based on the M51 galaxy from DustPedia. Different colors denote different observation angles, from $0^{\circ}$ (face-on) to $90^{\circ}$ (edge-on). An aperture radius of $20$ kpc was used. The galaxy dust density has been normalized so that the total optical depth along the R axis is $\tau_{V_{R}}=5$. The dashed black line reflects the true extinction $R_{V_{THEMIS}}$.
@@ 5 Conclusions
In this work we have used radiative transfer simulations to explore how SN Ia dust extinction and host galaxy dust attenuation relate to one another. More specifically, we have investigated under which conditions, if any, attenuation can be used as a proxy for extinction. Our main findings are the following:
@@
    SN Ia extinction (or “point source extinction” as used throughout this paper) and host galaxy attenuation are fundamentally different phenomena. When observing a galaxy, the light scattered into the line of sight becomes a very relevant component of the total observed flux, lowering the value of the effective $R_{V_{eff}}$ with respect to the medium’s intrinsic extinction $R_{V}$. In addition, sources at different depths contained along the same line of sight contribute to an increase in $R_{V_{eff}}$. When taken in conjunction, these two effects mean that the $R_{V_{eff}}$ measured for any given galaxy is fundamentally different from the optical extinction $R_{V}$ of its particular dust medium, even if small resolutions of $25$ pc are considered. These effects are further exacerbated if distinct stellar populations with different SEDs and spatial distributions are present in the galaxy, as is likely to occur. Thus, SN Ia standardization cannot be improved by using global or local $R_{V_{eff}}$ values obtained from SED stellar population fits to host galaxy data as proxies for the extinction of individual SNe. Doing so will only contribute to further bias the results, due to the geometric effects inherent to galaxy observations.
@@
    Global attenuation properties for the simulated galaxies are primarily driven by the observation angle. The global $A_{V_{eff}}-R_{V_{eff}}$ trends for observed galaxies, which have different morphologies, dust contents and stellar masses, can be qualitatively reproduced by a single simulated galaxy, with a single dust type, observed from different angles. This shows that the importance of geometric effects cannot be neglected in any analysis of galactic dust.
@@
    Galaxy dust density, as parameterized by the cloud normalization in the simulation, is directly linked to $A_{V_{eff}}$ and $R_{V_{eff}}$. Increasing the dust density leads to an increase in both quantities across all observation angles. As such, if a correlation between stellar mass and dust density is assumed, the geometric effects in the observations are also able to explain the observed increase in $R_{V_{eff}}$ values with stellar mass for star-forming spiral galaxies.
@@
    When spiral and elliptical galaxies are analyzed together, our simulations with a single dust type fail to explain some of the observed behaviors. The literature suggests that ellipticals have $A_{V_{eff}}$ values similar to low-mass spirals. However, if this is assumed in our model, the recovered $R_{V_{eff}}$ values, which are also supposed to be in line with those of low-mass spirals, are much higher than expected. This means that there is some intrinsic difference between the two galaxy types that cannot be modeled by geometrical and morphological effects, and that particularly relates to the optical properties of their specific dust types, with ellipticals having intrinsically lower $R_{V}$ values. In this regard, there could be multiple distinct dust populations associated with different galaxy populations that need to be corrected for accordingly in SN Ia standardization, as has previously been suggested in the literature.
@@
    All the geometrical effects contributing to $R_{V_{eff}}$ make it difficult to extract accurate information about intrinsic dust properties, even for well-resolved observations. However, our simulations suggest that, for elliptical and face-on spiral galaxies, one can recover $R_{V_{eff}}$ values within a $10\%$ error of the intrinsic extinction $R_{V}$ for sufficiently resolved areas with high $A_{V_{eff}}$. For edge-on spirals, the median $R_{V_{eff}}$ for low-$A_{V_{eff}}$ regions, although with a lot of scatter, appears to be better at tracing the intrinsic extinction $R_{V}$. This conclusion is tentatively supported by preliminary fits at resolved location of real observed galaxies, although it is difficult to confirm without direct knowledge of the intrinsic optical properties of the dust. In any case, it is crucial to ensure that the regions chosen to proxy extinction in SNe Ia or any other point sources allow for the elimination of both the scattered and the direct flux components coming from deeper regions of the galaxy. In this way, it might ultimately be possible to relate galaxy attenuation to the extinction of their hosted SNe Ia, which would allow individual SNe dust properties to be independently corrected. This would greatly improve upon SN Ia standardization, as it would mean forgoing the often assumed universal color-luminosity correction.
@@ Acknowledgements.
J. D., S. G. G. and A. M. acknowledge support by FCT for CENTRA through the Project No. UIDB/00099/2020. J. D. also acknowledges support by FCT under the PhD grant 2023.01333.BD, with DOI https://doi.org/10.54499/2023.01333.BD. S.G.G. acknowledges support from the ESO Scientific Visitor Programme. This work was funded by ANID, Millennium Science Initiative, ICN12_009. M. S. acknowledges support by the Ministry of Science, Technological Development and Innovation of the Republic of Serbia (MSTDIRS) through contract no. 451-03-136/2025-03/200002 with the Astronomical Observatory (Belgrade). Based in part on observations collected at the European Southern Observatory under ESO programme(s) 097.D-0408(A). Based in part on observations made with the NASA/ESA Hubble Space Telescope, obtained at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5-26555, associated with programs #17070 and #17191. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation. This work used public data from the GALEX Survey. GALEX is operated for NASA by the California Institute of Technology under NASA contract NAS5-98034. This publication makes use of data products from the Wide-field Infrared Survey Explorer, which is a joint project of the University of California, Los Angeles, and the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. This work makes use of simulations from DustPedia, a collaborative focused research project supported by the European Union under the Seventh Framework Programme (2007-2013) call (proposal no. 606847). The participating institutions are: Cardiff University, UK; National Observatory of Athens, Greece; Ghent University, Belgium; Université Paris Sud, France; National Institute for Astrophysics, Italy and CEA, France.